solve the initial value problem y''-4y'-5y=0 for y(0)=0 and y'(0)=6


1
Expert's answer
2021-09-07T10:15:50-0400

Auxiliary equation,

m2-4m-5=0

solving the equation, we get

m=-1, 5

Therefore, general solution is

"y=c_1e^{-x}+c_2e^{5x}"

Then,

"y'=-c_1e^{-x}+5c_2e^{5x}"

By using initial conditions,

"y(0)=c_1+c_2=0\\
and\\
y'(0)=-c_1+5c_2=6"

solving above two equations, we get

"c_1=-1 ,c_2=1"

Therefore, solution of the given initial value problem is y=-e-x +e5x.